An air-filled pipe is found to have successive harmonics at 800 Hz , 1120 Hz , and 1440 Hz . It is unknown whether harmonics below 800 Hz and above 1440 Hz exist in the pipe. What is the length of the pipe?

L= 53.6 cm

open-closed pipe

Need help with: Identify the correct displacement graph of the 1120 Hz standing wave in the pipe. Note that the closed end of the pipe is on the right.

An air-filled pipe is found to have successive har
Answer #1

Conreck answer option (a) #313 YnIs(velocity of s ima
Answer #2

It has to be an open closed pipe.

For an open closed pipe.

Let the fundamental frequency be f

So, the second harmonic = 3f

3rd harmonic = 5f

So, difference between harmonics = 3f - f = 2f

So, 2f = (1120-800) = 1440-1120 = 320

So, f = 160 Hz

So, 800Hz is the 800/160 = 5th harmonic

1120 Hz = 7th

1440 = 9th

Now, for an open - closed pipe,

f = v/4L

where v = speed of sound in air = 343 m/s

So, 160 = 343/4L

So, L = 0.536 m = 53.6 cm

A)

The unknowns are : Length of pipe , fundamental frequency, Mode numbers for given harmonics

The knowns are : Speed of sound